# Measuring quantum relative entropy with finite-size effect

### v2

@@ IV Analysis with increasing $d$
Now, we consider when the error goes to zero under the condition $d\to\infty$. We denote the $d$-dimensional system by ${\cal H}_{d}$. To discuss this case, we consider the sequence of states $\sigma_{d}$ depending on the dimension $d$. We also choose a subset ${\cal S}_{d}$ of the set ${\cal S}({\cal H}_{d})$ of states on ${\cal H}_{d}$. We assume the following condition for $\sigma_{d}$ and ${\cal S}_{d}$;
@@
$\leq$ $\|\log\rho-(\mathrm{Tr}\rho\log\rho)\|_{\rho}+\|\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}.$ (30)
@@
$\|\log\rho-(\mathrm{Tr}\rho\log\rho)\|_{\rho}^{2}=\mathrm{Tr}\rho(\log\rho-(\mathrm{Tr}\rho\log\rho))^{2}$
$=$ $(\mathrm{Tr}\rho(\log\rho)^{2})-(\mathrm{Tr}\rho\log\rho)^{2}\leq(\mathrm{Tr}\rho(\log\rho)^{2})\stackrel{{\scriptstyle(a)}}{{\leq}}(\log d)^{2},$ (31)
@@
$\|\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}^{2}=\mathrm{Tr}\rho(\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d}))^{2}$
$=$ $(\mathrm{Tr}\rho(\log\sigma_{d})^{2})-(\mathrm{Tr}\rho\log\sigma_{d})^{2}\leq(\mathrm{Tr}\rho(\log\sigma_{d})^{2})\leq(td)^{2}.$ (32)
@@
$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon}+\delta_{\rho,\sigma}^{(n),-,\epsilon}\leq$ $\frac{1}{\epsilon^{2}}\Big{(}\frac{\sqrt{c_{0}}}{\sqrt{c}}+\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}\Big{)}^{2}$
@@
Now, we compare the sample complexity of our method with the sample complexity of the simple application of full tomography [27, 34]. In particular, we employ the error analysis on the trance norm [27] rather than other error criteria because the trance norm is most useful to evaluate the error of the relative entropy. Their method achieves the estimation error
@@
We choose $s\in(0,1)$. Using Weyl’s dimension formula [32, Theorem 7.1.7], we have
@@
$=$ $\!\!\sum_{1\leq i<j\leq d}\!\!\log\Big{(}1+\frac{\lambda_{j}-\lambda_{i}}{j-i}\Big{)}=\sum_{l=1}^{d}\sum_{i=1}^{d-l}\log\Big{(}1+\frac{\lambda_{l+i}-\lambda_{i}}{l}\Big{)}$
@@
V Analysis on time complexity
@@
First, we apply the Schur transform $U_{Schur}$, which realizes a basis change to block-diagonalize $U^{\otimes n}$ for an arbitrary unitary $U$. The references [26, 36, 37, 38] proposed its quantum algorithm. We evaluate its prevision as follows. Given a required error $\epsilon>0$ and a target unitary $V$, we say that a unitary $U$ is an $\epsilon$-approximate of $V$ when any pure state $|\psi\rangle$ satisfies
@@
where $\bm{\lambda}!:=\lambda_{1}!\lambda_{2}!\cdots\lambda_{d}!$. For its detail, see [33, Eq. (2.72)]. Using [28, Theorem 2.5], we have
@@
$=$ $\sum_{j>i}\log(1+\frac{\lambda_{i}}{\lambda_{j}-\lambda_{i}+j-i})\leq\sum_{j>i}\log(1+\frac{\lambda_{i}}{j-i})$
@@
$\leq$ $\sum_{l=1}^{d}(d-l)\log\Big{(}1+\frac{n}{(d-l)l}\Big{)}\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}},$ (52)
@@
Given an arbitrary number $\epsilon>0$, we choose $c=\frac{16}{\epsilon}$. When $n=cd^{2}$, the limit of the upper bound in (53) can be bounded by $\epsilon$. Then, we define
@@
$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon,\epsilon_{1}}+\delta_{\rho,\sigma}^{(n),-,\epsilon,\epsilon_{1}}\leq\frac{(\sqrt{c_{0}}+4)^{2}}{c(\epsilon-\frac{4}{\sqrt{c}})^{2}}+\epsilon_{1}^{2}.$ (56)
@@
The paper [7] also mentions that there exists an algorithm for Schur transform $U_{Schur}$ whose time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$ when $n=O(d^{2})$ [39]. First, they employ the algorithm for weak Schur sampling given in [40, Section 4.2.2]. Although this method employs the quantum Fourier transform the permutation group $\mathfrak{S}_{n}$, they proposed to use the algorithm by [41] for this part. The obtained algorithm has time complexity $O(n^{3}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon))$. When $n=d^{2}$, it equals $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$. This implementation of weak Schur sampling was mentioned earlier in [42, Section 1.1.2].
VI Discussion
@@ Acknowledgement
The author is supported in part by the National Natural Science Foundation of China (Grant No. 62171212). The author is grateful for Professor Kun Fang to a helpful discussion. Also, the author would like to thank Mr Dmitry Grinko and Dr. Qisheng Wang for helpful discussions on the time complexity of Schur transform. In addition, the author would like to thank Prof. Jun Suzuki and Dr. Francesco Albarelli for helpful discussions on the Cramér-Rao type bound. Also, the author would like to thank Prof. Macro Tomamichel and Dr. Hu Yanglin for helpful discussions for the relation with full tomography and informing the references [27, 34].
@@ Appendix A Evaluation of mean square error: Proof of Theorem 2
$=$ $\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)\Big{(}\frac{1}{n}(-\log\langle u_{\bm{\lambda},j}|\sigma_{\bm{\lambda}}|u_{\bm{\lambda},j}\rangle-D(\rho\|\sigma)\Big{)}^{2}$
@@
$\Big{\|}-\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{\|}_{\rho^{\otimes n}}$
@@ Appendix B Local asymptotic normality: Proof of Theorem 3
We use the methods [3, 20, 44].
$=$ $-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle+\log d_{\bm{\lambda}}^{\prime})-D(\rho\|\sigma)$
The random variable $\sqrt{n}(-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma))$ has additional variable $j^{\prime}$. The joint distribution of $(\bm{\lambda},j,j^{\prime})$ is given as $\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$. This random variable coincides with $\sqrt{n}(\frac{1}{n}(\log{\cal E}_{\sigma^{\otimes n}}(\rho^{\otimes n})-\log\sigma^{\otimes n})-D(\rho\|\sigma))$ under the state ${\cal E}_{\sigma^{\otimes n}}(\rho^{\otimes n})$ studied in Section V of [20]. The discussion in Section V of [20] showed that it converges to Gaussian distribution with variance $V(\rho\|\sigma)$ in probability as $n\to\infty$.
@@ C.1 Preparation
We define the pinching map $\Gamma_{\rho,n}$ as
$\Gamma_{\sigma,n}(\rho^{\prime})$
::=:= $\sum_{\bm{\lambda}\in Y_{d}^{n},j}(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))\rho^{\prime}(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))$ (63)
for a general state $\rho^{\prime}$ on ${\cal H}^{\otimes n}$. Since the system ${\cal V}_{\bm{\lambda}}$ has no information, it is sufficient to handle the two states $\oplus_{\bm{\lambda}\in Y_{d}^{n}}\rho_{\bm{\lambda}}$ and $\oplus_{\bm{\lambda}\in Y_{d}^{n}}\sigma_{\bm{\lambda}}$ over $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$. We define the map
@@
$\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{\alpha}(\sigma^{\otimes n})^{1-\alpha}$
$=$ $\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\sigma,n}(\rho^{\otimes n})(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $\mathrm{Tr}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha}$
@@
$\cdot\big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}\big{)}$
$=$ $\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
$\cdot\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n,d}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
$\cdot((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
$=$ $d_{n,d}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$

### v4

@@ 4 Analysis with increasing $d$
We now discuss the behavior of $n$ with respect to $d$ that allows the error to converge to zero as $d\to\infty$, rather than assuming $n=O(d^{4+\delta})$. We denote the $d$-dimensional system by ${\cal H}_{d}$. To discuss this case, we consider the sequence of states $\sigma_{d}$ depending on the dimension $d$. We also choose a subset ${\cal S}_{d}$ of the set ${\cal S}({\cal H}_{d})$ of states on ${\cal H}_{d}$. We assume the following condition for $\sigma_{d}$ and ${\cal S}_{d}$;
@@
$\leq$ $\|\log\rho-(\mathrm{Tr}\rho\log\rho)\|_{\rho}$
$+\|\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}.$ (33)
@@
$\|\log\rho-(\mathrm{Tr}\rho\log\rho)\|_{\rho}^{2}$
$=$ $\mathrm{Tr}\rho(\log\rho-(\mathrm{Tr}\rho\log\rho))^{2}$
$=$ $(\mathrm{Tr}\rho(\log\rho)^{2})-(\mathrm{Tr}\rho\log\rho)^{2}$
$\leq$ $(\mathrm{Tr}\rho(\log\rho)^{2})\stackrel{{\scriptstyle(a)}}{{\leq}}(\log d)^{2},$ (34)
@@
$\|\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}^{2}$
$=$ $\mathrm{Tr}\rho(\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d}))^{2}$
$=$ $(\mathrm{Tr}\rho(\log\sigma_{d})^{2})-(\mathrm{Tr}\rho\log\sigma_{d})^{2}$
$\leq$ $(\mathrm{Tr}\rho(\log\sigma_{d})^{2})\leq(td)^{2}.$ (35)
@@
$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon}+\delta_{\rho,\sigma}^{(n),-,\epsilon}$
$\leq$ $\frac{1}{\epsilon^{2}}\Big{(}\frac{\sqrt{c_{0}}}{\sqrt{c}}+\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}\Big{)}^{2}$
@@
Now, we compare the sample complexity of our method with the sample complexity of the simple application of full tomography [33, 39]. In particular, we employ the error analysis on the trace norm [33] rather than other error criteria because the trace norm is most useful to evaluate the error of the relative entropy. Their method achieves the estimation error
@@
In fact, the paper [39] also considers the sample complexity when Bures $\chi^{2}$-divergence and Bures distance are adopted. However, the evaluation of the error under these criteria does not fit well with the above method. Moreover, it is challenging to find a good upper bound of the estimation errors of $S(\hat{\rho})-S(\rho)$ and $D(\hat{\rho}\|\sigma)-D(\rho\|\sigma)$ by using the estimation error based on these criteria.
@@
We choose $s\in(0,1)$. Using Weyl’s dimension formula [38, Theorem 7.1.7], we have
@@
$=$ $\!\!\sum_{1\leq i<j\leq d}\!\!\log\Big{(}1+\frac{\lambda_{j}-\lambda_{i}}{j-i}\Big{)}$
$=$ $\sum_{l=1}^{d}\sum_{i=1}^{d-l}\log\Big{(}1+\frac{\lambda_{l+i}-\lambda_{i}}{l}\Big{)}$
@@
5 Analysis on time complexity
@@
The references [32, 41, 42, 43] proposed quantum algorithms for the Schur transform $U_{Schur}$. We evaluate its precision as follows. Given a required error $\epsilon>0$ and a target unitary $V$, we say that a unitary $U$ is an $\epsilon$-approximate of $V$ when any pure state $|\psi\rangle$ satisfies
@@
where $\bm{\lambda}!:=\lambda_{1}!\lambda_{2}!\cdots\lambda_{d}!$. For its detail, see [37, Eq. (2.72)]. Using [19, Theorem 2.5], we have
@@
$=$ $\sum_{j>i}\log(1+\frac{\lambda_{i}}{\lambda_{j}-\lambda_{i}+j-i})$
$\leq$ $\sum_{j>i}\log(1+\frac{\lambda_{i}}{j-i})$
@@
$\leq$ $\sum_{l=1}^{d}(d-l)\log\Big{(}1+\frac{n}{(d-l)l}\Big{)}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}},$ (55)
@@
Given an arbitrary number $\epsilon>0$, we choose $c=\frac{16}{\epsilon^{2}}$. When $n=cd^{2}$, the limit of the upper bound in (56) can be bounded by $\epsilon$. Then, we define
@@
$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon,\epsilon_{1}}+\delta_{\rho,\sigma}^{(n),-,\epsilon,\epsilon_{1}}$
$\leq$ $\frac{(\sqrt{c_{0}}+4)^{2}}{c(\epsilon-\frac{4}{\sqrt{c}})^{2}}+\epsilon_{1}^{2}.$ (59)
@@
The paper [7] also mentions that there exists an algorithm for Schur transform $U_{Schur}$ whose time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$ when $n=O(d^{2})$ [44]. First, they employ the algorithm for weak Schur sampling given in [45, Section 4.2.2]. Although this method employs the quantum Fourier transform the permutation group $\mathfrak{S}_{n}$, they proposed to use the algorithm by [46] for this part. The obtained algorithm has time complexity $O(n^{3}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon))$. When $n=d^{2}$, it equals $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$. This implementation of weak Schur sampling was mentioned earlier in [47, Section 1.1.2].
Remark 9.
Recently, the paper [48] studied the estimation of relative entropy with two unknown densities $\rho$ and $\sigma$ when an estimator of the state is given and the estimator converges to the true density. The analysis [48] is different from ours in the following points. (i) We assume that $\sigma$ is known, but the paper [48] assumes that it is unknown. (ii) We discuss the optimality of our estimator in various settings. Under the case with fixed $d$, we optimize the MSE. When $d$ increases, we discuss the sample complexity. But, when an estimator of the state is given, the paper [48] discusses the behavior of the estimator of relative entropy by using the behavior of the estimator. In addition, the paper [48] mentioned that the limiting distribution of the normalized estimator is the normal distribution. When $d$ is fixed, the discussion explains how this problem is reduced to a special case of the parameter estimation with nuisance parameters. The paper [21] also showed that the limiting distribution of the normalized estimator is the normal distribution in the parameter estimation with nuisance parameters.
6 Discussion
@@ Acknowledgement
The author is supported in part by the National Natural Science Foundation of China (Grant No. 62171212) and the General R&D Projects of 1+1+1 CUHK-CUHK(SZ)-GDST Joint Collaboration Fund (Grant No. GRDP2025-022). The author is grateful for Professor Kun Fang to a helpful discussion. Also, the author would like to thank Mr Dmitry Grinko and Dr. Qisheng Wang for helpful discussions on the time complexity of Schur transform. In addition, the author would like to thank Prof. Jun Suzuki and Dr. Francesco Albarelli for helpful discussions on the Cramér-Rao type bound. Also, the author would like to thank Prof. Macro Tomamichel and Dr. Hu Yanglin for helpful discussions for the relation with full tomography and informing the references [33, 39].
@@ Appendix A Evaluation of mean square error: Proof of Theorem 2
We use the methods [3, 26, 49]. Then, we have the following calculation.
$=$ $-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle+\log d_{\bm{\lambda}}^{\prime})$
$-D(\rho\|\sigma)$
Therefore, we have
@@
$=$ $\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)\Big{(}\frac{1}{n}(-\log\langle u_{\bm{\lambda},j}|\sigma_{\bm{\lambda}}|u_{\bm{\lambda},j}\rangle$
$-D(\rho\|\sigma)\Big{)}^{2}$
@@
$\Big{\|}-\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{\|}_{\rho^{\otimes n}}^{2}$
@@ Appendix B Local asymptotic normality: Proof of Theorem 3
The random variable $\sqrt{n}(-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma))$ has additional variable $j^{\prime}$. The joint distribution of $(\bm{\lambda},j,j^{\prime})$ is given as $\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$.
We define the pinching map $\Gamma_{\sigma,n}$ as
$\Gamma_{\sigma,n}(\rho^{\prime})$
::=:= $\sum_{\bm{\lambda}\in Y_{d}^{n},j}(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))\rho^{\prime}$
for a general state $\rho^{\prime}$ on ${\cal H}^{\otimes n}$, where $P({\cal V}_{\bm{\lambda}})$ is the projection to ${\cal V}_{\bm{\lambda}}$. Then, this random variable coincides with $\sqrt{n}(\frac{1}{n}(\log\Gamma_{\sigma,n}(\rho^{\otimes n})-\log\sigma^{\otimes n})-D(\rho\|\sigma))$ under the state $\Gamma_{\sigma,n}(\rho^{\otimes n})$ studied in Section V of [26]. The discussion in Section V of [26] showed that it converges to Gaussian distribution with variance $V(\rho\|\sigma)$ in probability as $n\to\infty$.
@@ C.1 Preparation
Since the system ${\cal V}_{\bm{\lambda}}$ has no information, it is sufficient to handle the two states $\oplus_{\bm{\lambda}\in Y_{d}^{n}}\rho_{\bm{\lambda}}$ and $\oplus_{\bm{\lambda}\in Y_{d}^{n}}\sigma_{\bm{\lambda}}$ over $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$. We define the map
@@
$\mathrm{Tr}\big{(}\Gamma_{\sigma,n}(\rho^{\otimes n})\big{)}^{\alpha}(\sigma^{\otimes n})^{1-\alpha}$
$=$ $\mathrm{Tr}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\sigma,n}(\rho^{\otimes n})(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha}$
$=$ $\mathrm{Tr}\Big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}^{\alpha}$
@@
$\cdot\Big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}^{\alpha-1}$
$=$ $\mathrm{Tr}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$\cdot\Big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}^{\alpha-1}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n,d}^{1-\alpha}\mathrm{Tr}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$\cdot\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}$
$=$ $d_{n,d}^{1-\alpha}\mathrm{Tr}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha}$